The differential equation of damped vibrating system under the action of an external periodic force is d^2x/dt^2+ 2r(dx/dt)+ n^2x= acospt
1
Expert's answer
2020-10-27T17:10:21-0400

"\displaystyle
\frac{\mathrm{d}^2x}{\mathrm{d}t^2} + 2r\left(\frac{\mathrm{d}x}{\mathrm{d}t}\right)+ n^2x= a\cos(pt) \\

\textsf{Let}\, x = f(t) \\

f"(t) + 2rf'(t) + n^2f(t) = a\\cos(pt)\\\\\n\n\\textsf{The complementary function is derived}\\\\\n\\textsf{when}\\, \nf"(t) + 2rf'(t) + n^2f(t) = 0\\

\text{The auxiliary equation is}\\


m^2 + 2rm + n^2 = 0\\


m = \frac{-2r \pm \sqrt{4r^2 - 4n^2}}{2} = -r \pm \sqrt{r^2 - n^2}\\

\textsf{Since}\, n > r\\

m = -r \pm j\sqrt{n^2 - r^2}\\

\textsf{Recall that if the auxiliary equation is}\\
m = \alpha \pm j\beta, \, \textsf{the general solution can be}\\
\textsf{written as}\\

y = e^{\alpha t}(A\cos(\beta t) + B\sin(\beta t))\\

\therefore \textsf{The complementary function of the given DE is}\\

f(t) = e^{-rt}(C\cos(\sqrt{n^2 - r^2}t + D\sin(\sqrt{n^2 - r^2}))\\


\textsf{This is a system with a frequency of}\\ \frac{2\pi}{\sqrt{n^2 - r^2}}\, \textsf{per units of time and a}\\
\textsf{period of}\, \sqrt{n^2 - r^2} \, \textsf{units of time and an amplitude}\\
\textsf{that is decreasing with time.}\\

\textsf{The effect of the}\, f'(t)\, \textsf{in the differential}\\
\textsf{equation is to introduce damping}\\\textsf{into the oscillatory motion so}\\\textsf{causing it to decay.}\\


\textsf{The coefficient}\, e^{-rt} \, \textsf{introduced damping}\\
\textsf{into the system, and since}\\
\textsf{as}\, t \rightarrow \infty, e^{-rt} \rightarrow 0 \implies f(t)\rightarrow 0,\,\textsf{the vibrations}\\
\textsf{are ultimately damped out.}\\


\textsf{The particular integral is of the form}\\

f(t) = A\cos(pt) + B\sin(pt)\\

f'(t) = -Ap\sin(pt) + Bp\cos(pt)\\


f''(t) = -Ap^2\cos(pt) - Bp^2\sin(pt)\\


 -Ap^2\cos(pt) - Bp^2\sin(pt) + 2r(-Ap\sin(pt) + Bp\cos(pt)) + n^2(A\cos(pt) + B\sin(pt)) = a\cos(pt)\\


(-Ap^2 + 2rBp + An^2)\cos(pt) + (-Bp^2 - 2rAp + Bn^2)\sin(pt) = a\cos(pt)\\

\textsf{Comparing coefficients}\\

-Ap^2 + 2rBp + An^2 = a\\

-Bp^2 - 2rAp + Bn^2 = 0\\

A(n^2 - p^2) + 2rBp = a\hspace{0.5cm}(1)\\

B(n^2 - p^2) - 2rAp = 0\hspace{0.5cm}(2)\\

(1) \times B \, \textsf{yields}\\
AB(n^2 - p^2) + 2rB^2p = aB\hspace{0.5cm}(3)\\

(2) \times A \, \textsf{yields}\\
AB(n^2 - p^2) - 2rA^2p = 0\hspace{0.5cm}(4)\\

\textsf{Subtract}\, (4) \, \textsf{from}\, (3)\\

2rA^2 + 2rB^2 = aB\\


A^2(n^2 - p^2) + 2rABp = aA\hspace{0.5cm}(5)\\

B^2(n^2 - p^2) - 2rABp = 0\hspace{0.5cm}(6)\\

\textsf{Add}\, (5)\, \textsf{and}\, (6)\\

(A^2 + B^2)(n^2 - p^2) = aA\hspace{0.5cm}(7)\\

2r(A^2 + B^2) = aB\hspace{0.5cm}(8)\\

\implies A^2 + B^2 = \frac{aB}{2r}\\

\textsf{Inserting in}\, (7)\\

\frac{aB}{2r}\cdot(n^2 - p^2) = aA\\

B(n^2 - p^2) = 2rA\\

\implies B = \frac{2rA}{n^2 - p^2}\\
\textsf{By}\, (5) \,\textsf{and} \, (6)\\

A + \frac{2rpB}{n^2 - p^2} = \frac{a}{n^2 - p^2}\\

-\frac{2rpA}{n^2 - p^2} + B = 0\\

-\frac{4rp^2 A}{(n^2 - p^2)^2} + \frac{2rpB}{n^2 - p^2} = 0\\


A\left(1 + \frac{(2rp)^2}{(n^2 - p^2)}\right) = \frac{a}{n^2 - p^2}\\

A\left(\frac{(n^2 - p^2)^2 + 4r^2p^2}{(n^2 - p^2)^2}\right) = a\\


A = \frac{a(n^2 - p^2)}{(n^2 - p^2)^2 + 4r^2p^2}\\

B =\frac{2r}{n^2 - p^2}\\ A = \frac{2arp}{(n^2 - p^2)^2 + 4r^2p^2}\\

\therefore f(t) = \frac{a(n^2 - p^2)}{(n^2 - p^2)^2 + 4r^2p^2}\cos(pt) + \frac{2arp}{(n^2 - p^2)^2 + 4r^2p^2}\sin(pt)\\
\begin{aligned}
f(t) &= \frac{a}{(n^2 - p^2)^2 + 4r^2p^2}\sqrt{(2rp)^2 + (n^2 - p^2)}\cos\left(pt - \arctan\left(\frac{2r}{n^2 - p^2}\right)\right)\\
&=\frac{a}{\sqrt{(n^2 - p^2)^2 + 4r^2p^2}}\cos\left(pt - \arctan\left(\frac{2r}{n^2 - p^2}\right)\right)
\end{aligned}\\
\textsf{Thus, the Particular integral is of the form}\\ b\cos(pt - \alpha),\, \textsf{where}\, b^2 = \frac{a^2}{(n^2 - p^2)^2 + 4r^2p^2}\\


\textsf{And the solution to the differential equation is}\\

x = e^{-rt}(C\cos(\sqrt{n^2 - r^2}t + D\sin(\sqrt{n^2 - r^2})) + \frac{a}{\sqrt{(n^2 - p^2)^2 + 4r^2p^2}}\cos\left(pt - \arctan\left(\frac{2r}{n^2 - p^2}\right)\right)\\"